Need assistance please.
The width of a rectangle is 6 feet shorter than its length. If its perimeter is 82 feet, what is its
width?
I set it up as follows:
82=2(x-6)+2x
Which gave me a solution of
47/2=x
23.5=x
I cannot figure out what to do next.
Thank you

Answer by ikleyn(52786)

.
In your notations, x is the length (although it would be more logical to make it the width . . . )


82 = 2(x-6) + 2x   <<<---=== Correct !


Next steps should be:

41 = x - 6 + x


41 + 6 = 2x


47 = 2x


x = 47/2 = 23.5    <<<---=== it is your LENGTH


Then you should find the WIDTH:  W = 23.5 - 6 = 17.5.


The problem is just solved and the solution is completed.